# Generative Software Engineering

### v1

@@ Abstract.
The rapid development of deep learning techniques, improved computational power, and the availability of vast training data have led to significant advancements in pre-trained models and large language models (LLMs). Pre-trained models based on architectures such as BERT and Transformer, as well as LLMs like ChatGPT, have demonstrated remarkable language capabilities and found applications in Software engineering. Software engineering tasks can be divided into many categories, among which generative tasks are the most concern by researchers, where pre-trained models and LLMs possess powerful language representation and contextual awareness capabilities, enabling them to leverage diverse training data and adapt to generative tasks through fine-tuning, transfer learning, and prompt engineering. These advantages make them effective tools in generative tasks and have demonstrated excellent performance. However, there are limitations in the existing literature reviews on the development of pre-trained models and LLMs in the context of software engineering: (1) either only focus on the pre-trained models or only on LLMs, lacking the analysis of the development process from the pre-trained models to LLMs. (2) Software engineering tasks are not distinguished by type. (3) Lack of systematic analysis. In this paper, we present a comprehensive literature review of generative tasks in SE using pre-trained models and LLMs. We accurately categorize SE generative tasks based on software engineering methodologies and summarize the advanced pre-trained models and LLMs involved, as well as the datasets and evaluation metrics used. Additionally, we identify key strengths, weaknesses, and gaps in existing approaches, and propose potential research directions. This review aims to provide researchers and practitioners with an in-depth analysis and guidance on the application of pre-trained models and LLMs in generative tasks within SE.
@@ 1. Introduction
In the field of natural language processing, with the rapid development of deep learning techniques, improved computational power, and the continuous expansion of training data, mainstream approaches have undergone significant transitions from traditional models to pre-trained models and then to large language models (LLMs). Pre-trained models based on architectures such as BERT and Transformer, as well as LLMs represented by ChatGPT, have access to vast and diverse training data, showcasing powerful language capabilities comparable to humans and being applied in various domains. With the ability to learn from and generate text from large-scale corpora, powerful pre-trained models and LLMs are closing the gap between natural language and machine language.
@@
Software engineering is a discipline that focuses on the development, implementation, and maintenance of software systems. It is also one of the important application areas for pre-trained models and LLMs. Downstream tasks in software engineering can be categorized into two types: classification and generation, which are also the two major types of natural language processing tasks. Pre-trained models and LLMs possess powerful language representation and contextual awareness capabilities, enabling them to leverage diverse training data and adapt to generative tasks through fine-tuning, transfer learning, and prompt engineering. These advantages make them effective tools in generative tasks and have demonstrated excellent performance in the field of software engineering. There have been some literature reviews on the relationship between SE (Software Engineering) and pre-trained models and LLM (Language Models). However, these studies have some limitations: (1) Some studies narrowly focus on either pre-trained models or LLMs, ignoring the development stages from pre-trained models to LLMs. For example, Mosel et.al. (von der Mosel et al., 2023) focus on the ability of Transformer-based pre-trained models to understand words and sentences in the context of software engineering, while Hou et.al. (Hou et al., 2023) lean towards LLM4SE, particularly emphasizing how to use LLM to optimize processes and results. (2) Some studies mix different types of SE tasks, failing to distinguish between the two major types of natural language processing tasks. For example, although Zheng et.al. (Zheng et al., 2023a) divides software engineering tasks into seven types, they do not classify these tasks into classification tasks or generation tasks. (3) Alternatively, some studies only explore the performance of LLM in various SE tasks through empirical experiments, without conducting systematic literature reviews. At the same time, considering the application of LLM in SE generation tasks from pre-trained models is a complex task. It requires considering representative models at different stages of development, understanding the unique features of different models, considering fine-tuning strategies of pre-trained models and prompt engineering of LLM, conducting thorough literature reviews and data analysis, and overcoming challenges in the implementation process. Currently, there is a lack of detailed scrutiny and review of pre-trained models and LLM specifically for generative tasks in software engineering in the existing literature. Therefore, our research aims to fill this gap, analyze and summarize the performance of pre-trained models and LLM in SE generation tasks, and provide development guidance in this direction.
@@
In this paper, we provide a systematic literature review of generative tasks in Software Engineering (SE) based on pre-trained models and Large Language Models (LLM). By summarizing and comparing state-of-the-art technologies and the developmental history of models, we identify the key strengths, weaknesses, and gaps in existing generative task methods. Additionally, we propose potential directions for future research. Our review aims to provide researchers and practitioners with a comprehensive analysis and guide on the application of pre-trained models and LLM in generative tasks within SE. We anticipate that our findings will contribute to guiding future work in this rapidly evolving field. Our key contributions are as follows:
@@ 2. Methodology
Table 1. Retrieval keywords and statistics about generative tasks.
@@
Generation Large
Language Model,
@@
LLM,
Requirements
Large
Language
Model
Code
Generation
LLM, Code
@@
Code LLM,
Code Large
@@
LLM,
Test Oracle
@@
LLM,
Test Assert
Generation
LLM
Patch
Generation,
@@
LLM, Code
@@
LLM, Code
@@
  • The research date should be after [date].
To ensure the accuracy and efficiency of the literature screening process, we employed a closed card sorting method (Kitchenham and Charters, 2007) to categorize the collected literature quickly based on the above criteria into relevant and irrelevant papers. Specifically, for each generative subtask, the researcher responsible for literature collection applied the closed card sorting method and screening criteria to the papers returned by the search engine. In cases where the inclusion of a paper was ambiguous, these papers were compiled separately, and discussions were held to decide whether they met our requirements. Through this process, we ultimately identified [number] papers that met our research criteria.
@@
With the rapid development of LLM, an increasing number of researchers have started investigating its performance in software engineering tasks. Code generation, being the most popular downstream task in software engineering, has garnered significant attention, leading to numerous publications on evaluation datasets, methods, and related papers. The research on code summarization generation, test case generation, and patch generation is moderate, while there is relatively less research on tasks like requirements generation and code optimization, which only amounts to [number] publications.
@@ 2.1.1. Datasets
Requirements dataset (user stories)(F.Dalpiaz, [n. d.]). Dalpiaz et al.(F.Dalpiaz, [n. d.]) proposed a dataset comprising 22 product backlogs and 1,679 user stories meticulously curated by the requirements engineering community. Published as a raw archive, it consists of 22 text files, each dedicated to a specific product and containing associated user stories, presented in a single line per story. To the best of our knowledge, there is currently no publicly accessible expert-based annotation that establishes a definitive reference for the various concepts described within this dataset.
@@
Ace-design/qualified-user-stories dataset(S. Arulmohan and Meurs, 2023).The dataset created by Arulmohan et al.(Arulmohan et al., 2023) comprises an annotated version of Requirements dataset(user stories)(F.Dalpiaz, [n. d.]), encompassing 22 products and 1,679 user stories. They first started by manually annotating the Requirements dataset using the Doccano platform [^1] https://doccano.herokuapp.com. They loaded the labels (i.e., Persona, Action, Entity, Benefit) and two relations (i.e., triggers, targets) during their processing. To complement the ground truth, they ran VN(Robeer et al., 2016) on top of the provided backlogs and extracted the concepts by parsing the output files it produced. After this step, they obtain a new annotated dataset containing the results of a domain model extraction on top of the input corpus.
@@
Improved version of the PROMISE NFR datasethey2020norbert. Hey et al. (Hey et al., 2020) gathered 249 non-functional requirements from the PROMISE NFR dataset (J. Cleland-Huang and Port, 2007), spanning 15 projects. Each requirement statement is categorized into one of four NFR classes: usability, security, operational, and performance.
@@
App review NFR multi-Label classification dataset(Jha and Mahmoud, 2019). Jha et al. (Jha and Mahmoud, 2019) present a dataset comprising 1800 reviews spanning various categories from both Google Play and the Apple App Store. Each review is tagged with at least one NFR label, including dependability, performance, usability, supportability, or a miscellaneous label indicating that the review does not mention NFR.
2.1.2. Methods
Study of using LLMs technology to improve requirements engineering.White et al.(White et al., 2023) proposed a heuristic design technique for software engineering, presented in patterns, to address common challenges encountered when automating requirement engineering using LLMs such as ChatGPT. They explored several heuristic patterns that have already been applied to enhance requirements gathering in three different classifications: Requirements Simulator, Specification Disambiguation, and Change Request Simulation(example in Fig. 1). The patterns have already been applied to enhance requirements gathering. Their work also contributes to the research on applying LLMs in requirements engineering.
Figure 1. Examples of requirement simulator pattern
Wang et al.(Wang et al., 2023c) proposed a dialogue framework called Chatcoder between a large language model (LLM) and users. Within the framework, the LLM analyzes arguments to refine the user’s original requirement expression, then returns the refined arguments to users in an understandable format.The overall structure of ChatCoder comprises a two-round dialogue (Fig. 2). The first round, "Paraphrase and Extend," addresses vague or incomplete user requirements. ChatCoder prompts the LLM to paraphrase the user’s original requirements from different angles for clarity. For missing or ambitious arguments requiring refinement, ChatCoder asks the LLM to extend them using assumptions from its training data. Users review and correct the refined specifications.The second round, "Going-deep and Loop-back," involves the LLM asking users about confusion in refined specifications for further refinement. Users answer questions and correct specifications if based on incorrect requirements. After these rounds, the refined requirement is sent to LLMs to generate the desired programs.
Figure 2. An example of Chatcoder
Arulmohan et al.(Arulmohan et al., 2023) propose to investigate how LLMs and conversational agents can support model extraction from requirements, focusing on agile backlogs. Specifically, the process begins by setting up the system role, where they instruct the engine to adopt a persona and outline the overarching task for later request execution context setup. Moving on to the next phase, under the continued guise of the system role, they elaborate on extracting personas, entities, actions, and benefits from a story. Subsequently, they present an illustrative example of such extraction using a manually annotated story. Transitioning to the user role, they then supply the <STORY> to be processed. Given the statelessness of the model, the answer from the previous phase needs to be injected into the conversation. Consequently, adopting the assistant role, they introduce a conversation entry detailing the <CONCEPTS> acquired earlier. Following a similar pattern as in the preceding phase, they articulate the task using the system role (categorizing primary and secondary actions and entities) and provide an example of such an execution. The final step mirrors the previous patterns. They inject the <CATEGORIES> as the assistant, followed by task description and example provision in the system role. The conversation they used is described in Fig. 3.
Figure 3. An example of work process for extracting domain model
Zhang et al.(Zhang et al., [n. d.]) conducted a comprehensive evaluation of the performance of ChatGPT in the task of requirement information retrieval. This assessment offers profound insights for the design of more efficient retrieval methods. The evaluation results indicate that ChatGPT successfully retrieves relevant requirement information. However, under zero-shot conditions, its precision in retrieving specific requirement information is subject to certain limitations. These findings provide valuable guidance for further optimizing the application of ChatGPT in the field of requirement information retrieval.
Waseem et al.(Waseem et al., 2023) provide an in-depth exploration of the effects of students utilizing ChatGPT at different stages of the software development life cycle. Specifically, they provided students with a software project topic for development. During the requirements analysis phase, students gathered functional and non-functional requirements with ChatGPT to understand project needs comprehensively. The assistance of ChatGPT is crucial for subsequent development stages. Upon completion of the development, Waseem et al.(Waseem et al., 2023) surveyed to evaluate the effectiveness of ChatGPT, revealing a significant reduction in the time required for defining requirements during the requirements analysis phase. Additionally, students responded positively to the utility of ChatGPT, believing that its usage enhances requirement clarity. Also, ChatGPT aids in a better understanding of project needs and facilitates stakeholder communication. To conclude, ChatGPT demonstrated a significant positive impact during the requirements analysis phase of software development projects.
@@ 2.2.1. Datasets
Chen et.al. (Chen et al., 2021) utilized a purely handwritten collection of programming problems called HumanEval [^2] https://www.github.com/openai/human-eval. This dataset is based on the Python programming language and consists of 164 programming problems. A sample test case from HumanEval is illustrated in Figure 4. Each programming problem in HumanEval consists of a function name, a natural language statement describing the function of the function, and unit test cases used to evaluate the generated code. The average number of unit test cases for each programming problem is 7.1. These programming problems cover many types, including language comprehension, reasoning, algorithms, and simple mathematics.
Figure 4. An example of HumanEval
To assess the ability of LLM to generate concise Python code, MBPP [^3] https://github.com/google-research/google-research/tree/master/mbpp. (Austin et al., 2021) (The Mostly Basic Programming Problems) is proposed. The dataset includes 974 short Python programming problems, and the difficulty of each programming problem is set at the entry-level programmer level. Among these problems, 58% are mathematical calculation, 43% involve list processing, 19% require string operations, 9% process integer sequences, and 2% are centered on using other data structures. An illustrative sample of this dataset is shown in Figure 5, similar to HumanEval, each programming problem sample consists of a function name, the natural language describing the target code, and three unit test cases used to evaluate the generated code.
Figure 5. An example of MBPP
@@
Wang et.al. (Wang et al., 2023a) proposed a dataset called MCoNaLa [^6] https://github.com/zorazrw/multilingual-conala. with different natural languages as prompts. They annotated a total of 896 natural language-code pairs in three languages: Spanish, Japanese, and Russian, as illustrated in Figure 6. Specifically, they engaged annotators fluent in both the respective natural language and Python. The annotators manually collected 341, 210, and 345 natural language-code pairs from the Spanish, Japanese, and Russian subforums of StackOverflow, respectively.
Figure 6. Examples of MCoNaLa
@@
To further investigate the performance of LLMs in generating more complex code, Du et.al. (Du et al., 2023) proposed a class-level code generation benchmark dataset. They manually build a benchmark for 100 class-level Python code generation tasks, named ClassEval [^7] https://github.com/FudanSELab/ClassEval.. Each coding task in ClassEval comprises a natural language description of the target class, a test suite used to validate the correctness of the generated code, and a specification solution serving as a reference implementation for the target class. Typically, LLMs generate code snippets based on the input description and use the provided test suite to verify correctness. They build test cases on two levels: method-level testing and class-level testing. To ensure that the generated code can be effectively checked according to the given test suite, they designed class-level (import statement, class name, class description, and class constructor) and method-level (method signature, function description, parameter/return description, and sample input/output) descriptions for the target code. Method-level testing mainly involves independently calling each method to verify its correctness without calling any other method in the class. Class-level testing mainly evaluates correctness through sequential calls. Figure 7 illustrates an example of ClassEval.
Figure 7. Examples of ClassEval
Although the LLM code generation model has achieved exciting performance, it is often fragile, as slight modifications to prompts can result in significantly different outputs. Most existing work on robustness in text or code tasks has focused on classification, lacking benchmarks for testing the robustness of code generation. Therefore, Wang et.al. (Wang et al., 2023d) proposes ReCode, a comprehensive benchmark for evaluating the robustness of code generation models. They tailored over 30 perturbations specifically for code, including document strings, function and variable names, code syntax, and code formatting, as shown in Table 3. These perturbations were applied to the HumanEval and MBPP datasets, forming a robustness benchmark test dataset for code generation.
Table 3. The perturbations used by ReCode
On which Doc strings Function names Code Syntax Code Format
Perturbations Back Translation Camel Case Dead Code Inserter Newline Insertion
Butter Fingers Butter Fingers For-While Switch Tap-Indent
Change Char Case Swap Characters Operand Swap Line Split
English Inflectional Variation Change Char Case Var Renamer CB Doc strings to Comments
Swap Characters Inflectional Variation Var Renamer Naive
Synonym Insertion Synonym Substition Var Renamer RN
Synonym Substitution
Tense Transformation Past
Tense Transformation Future
Whitespace

### v2

@@ Abstract.
The rapid development of deep learning techniques, improved computational power, and the availability of vast training data have led to significant advancements in pre-trained models and large language models (LLMs). Pre-trained models based on architectures such as BERT and Transformer, as well as LLMs like ChatGPT, have demonstrated remarkable language capabilities and found applications in Software engineering. Software engineering tasks can be divided into many categories, among which generative tasks are the most concern by researchers, where pre-trained models and LLMs possess powerful language representation and contextual awareness capabilities, enabling them to leverage diverse training data and adapt to generative tasks through fine-tuning, transfer learning, and prompt engineering. These advantages make them effective tools in generative tasks and have demonstrated excellent performance. However, there are limitations in the existing literature reviews on the development of pre-trained models and LLMs in the context of software engineering: (1) either only focus on the pre-trained models or only on LLMs, lacking the analysis of the development process from the pre-trained models to LLMs. (2) Software engineering tasks are not distinguished by type. (3) Lack of systematic analysis. In this paper, we present a comprehensive literature review of generative tasks in SE by applying pre-trained models and LLMs. We categorize software engineering generative tasks based on the software development life cycle and summarize the advanced pre-trained models and LLMs involved, as well as the datasets and evaluation metrics used. Additionally, we identify key strengths, weaknesses, and gaps in existing approaches, and propose potential research directions. This review aims to provide researchers and practitioners with an in-depth analysis and guidance on the application of pre-trained models and LLMs in generative tasks within SE.
@@ 1. Introduction
Pre-trained models such as BERT (Devlin et al., 2019), Transformer (Vaswani et al., 2023), and LLMs, leveraging vast and diverse datasets, showcase formidable linguistic capabilities and find widespread applications across various domains, effectively bridging the gap between natural and machine languages(von der Mosel et al., 2023; Niu et al., 2022a). Pre-trained models based on architectures such as BERT and Transformer, as well as LLMs represented by ChatGPT, have access to vast and diverse training data, showcasing powerful language capabilities comparable to humans and being applied in various domains. With the ability to learn from and generate text from large-scale corpora, powerful pre-trained models and LLMs perform well on various generation tasks (Tang et al., 2023).
@@
Software engineering is a discipline that focuses on the development, implementation, and maintenance of software systems (Zan et al., 2023; Kotti et al., 2023). It is also one of the important application areas for pre-trained models and LLMs. Downstream tasks in software engineering can be broadly categorized into two types: classification and generation, which are also the two major types of natural language processing tasks. Pre-trained models and LLMs possess powerful language representation and contextual awareness capabilities, enabling them to leverage diverse training data and adapt to generative tasks through fine-tuning (Ding et al., 2023), transfer learning (Tan et al., 2018), and prompt engineering (Brown et al., 2020). These advantages make them effective tools in generative tasks and have demonstrated excellent performance in the field of software engineering. There have been some literature reviews on the relationship between SE (Software Engineering) and pre-trained models and LLM (Language Models). However, these studies have some limitations: (1) Some studies narrowly focus on either pre-trained models or LLMs, ignoring the development stages from pre-trained models to LLMs. For example, Mosel et.al. (von der Mosel et al., 2023) focus on the ability of Transformer-based pre-trained models to understand words and sentences in the context of software engineering, while Hou et.al. (Hou et al., 2023) lean towards LLM4SE, particularly emphasizing how to use LLM to optimize processes and results. (2) Some studies mix different types of SE tasks, failing to distinguish between the two major types of natural language processing tasks. For example, although Zheng et.al. (Zheng et al., 2023a) divides software engineering tasks into seven types, they do not classify these tasks into classification tasks or generation tasks. (3) Alternatively, some studies only explore the performance of LLM in various SE tasks through empirical experiments, without conducting systematic literature reviews. At the same time, considering the application of LLM in SE generation tasks from pre-trained models is a complex task. It requires considering representative models at different stages of development, understanding the unique features of different models, considering fine-tuning strategies of pre-trained models and prompt engineering of LLM, conducting thorough literature reviews and data analysis, and overcoming challenges in the implementation process. Currently, there is a lack of detailed scrutiny and review of pre-trained models and LLM specifically for generative tasks in software engineering in the existing literature. Therefore, our research aims to analyze and summarize the performance of pre-trained models and LLM in SE generation tasks, and provide development guidance in this direction.
@@
In this paper, we provide a systematic literature review of generative tasks in software engineering (SE) based on pre-trained models and Large Language Models (LLM). By summarizing and comparing state-of-the-art technologies and the developmental history of models, we identify the key strengths, weaknesses, and gaps in existing generative task. Additionally, we propose potential directions for future research. Our review aims to provide researchers and practitioners with a comprehensive analysis and guide on the application of pre-trained models and LLM in generative tasks within SE. Our key contributions are as follows:
@@ 2. Methodology
Table 1. Retrieval keywords and statistics about SE generative tasks using Pre-trained Models and LLMs.
@@
Generation
Pre-trained
Model,
@@
Requirements
LLM
Code
Generation
LLM,
Code
Generation
Pre-trained
Model, Code
Generation
Large
@@
Pre-trained
Model,
Test Case
@@
Large
Language
Model
@@
LLM,
Code
Optimization
Pre-trained
Model,
Code
@@
LLM,
Code
Translation
Pre-trained
Model,
Code
@@
  • The research date should be after 2017.
To ensure the accuracy and efficiency of the literature screening process, we employed a closed card sorting method (Kitchenham and Charters, 2007) to categorize the collected literature quickly based on the above criteria into relevant and irrelevant papers. Specifically, for each generative subtask, the researcher responsible for literature collection applied the closed card sorting method and screening criteria to the papers returned by the search engine. In cases where the inclusion of a paper was ambiguous, these papers were compiled separately, and discussions were held to decide whether they met our requirements. Through this process, we ultimately identified 91 papers that met our research criteria.
@@
With the rapid development of LLM, an increasing number of researchers have started investigating its performance in software engineering tasks. Code generation, being the most popular downstream task in software engineering, has garnered significant attention, leading to numerous publications on evaluation datasets, methods, and related papers. The research on code summarization generation, test case generation, and patch generation is moderate, while there is relatively less research on tasks like requirements generation and code optimization, which only amounts to 6 and 3 publications, respectively.
@@ 2.1.1. Datasets
Requirements dataset (user stories)(F.Dalpiaz, [n. d.]). Dalpiaz et al. (F.Dalpiaz, [n. d.]) proposes a dataset comprising 22 product backlogs and 1,679 user stories meticulously curated by the requirements engineering community. The dataset is published as a raw archive and includes 22 text files, each of which is dedicated to a specific product and contains associated user stories. Each story is presented in a single line. As far as we know, there is currently no publicly accessible expert-based annotation that establishes a definitive reference for the various concepts described within this dataset.
@@
Ace-design/qualified-user-stories dataset (S. Arulmohan and Meurs, 2023).The dataset created by Arulmohan et al. (Arulmohan et al., 2023) comprises an annotated version of Requirements dataset(user stories) (F.Dalpiaz, [n. d.]), encompassing 22 products and 1,679 user stories. They first start by manually annotating the Requirements dataset using the Doccano platform  [^1] https://doccano.herokuapp.com. They load the labels (i.e., Persona, Action, Entity, Benefit) and two relations (i.e., triggers, targets) during their processing. To complement the ground truth, they run VN (Robeer et al., 2016) on top of the provided backlogs and extract the concepts by parsing the output files it produced. After this step, they obtain a new annotated dataset containing the results of a domain model extraction on top of the input corpus.
@@
Improved version of the PROMISE NFR dataset (Hey et al., 2020). Hey et al. (Hey et al., 2020) gathers 249 non-functional requirements from the PROMISE NFR dataset (J. Cleland-Huang and Port, 2007), spanning 15 projects. Each requirement statement is categorized into one of four NFR classes: usability, security, operational, and performance.
@@
App review NFR multi-Label classification dataset (Jha and Mahmoud, 2019). Jha et al. (Jha and Mahmoud, 2019) presents a dataset comprising 1800 reviews spanning various categories from both Google Play and the Apple App Store. Each review is tagged with at least one NFR label, including dependability, performance, usability, supportability, or a miscellaneous label indicating that the review does not mention NFR.
@@ 2.1.2. Evaluation Criteria
2.1.3. Methods
Study of using LLMs technology to improve requirements engineering.White et al. (White et al., 2023) proposes a heuristic design technique for software engineering, presented in patterns, to address common challenges encountered when automating requirement engineering using LLMs such as ChatGPT. They explored several heuristic patterns that have already been applied to enhance requirements gathering in three different classifications: Requirements Simulator, Specification Disambiguation, and Change Request Simulation(example in Fig. 1). The patterns have already been applied to enhance requirements gathering. Their work also contributes to the research on applying LLMs in requirements engineering.
Figure 1. Examples of requirement simulator pattern
Wang et al. (Wang et al., 2023c) proposes a dialogue framework called Chatcoder between LLMs and users. Within the framework, LLMs analyze arguments to refine the user’s original requirement expression, then return the refined arguments to users in an understandable format. The overall structure of ChatCoder comprises a two-round phase (Fig. 2). In the first phase, ChatCoder is tasked to rewrite the user’s initial demand expression using a preset perspective extracted from existing requirements engineering research. Specifically, ChatCoder packages instructions for rewriting user requirements and preset angles for extension in a prompt to instruct large language models to perform rewriting. The prompt is then sent to the large language model to get response. During the second phase, the chatcoder asks the large language model to further refine the requirements based on the specifications obtained in the first phase. Then, the user is asked to review the refined requirements and check for any possible errors. Finally, the refined requirements are sent to the LLMs to generate the desired programs.
Figure 2. An example of Chatcoder
Arulmohan et al. (Arulmohan et al., 2023) proposes to investigate how LLMs and conversational agents can support model extraction from requirements, focusing on agile backlogs. Specifically, the process begins by setting up the system role, where they instruct the engine to adopt a persona and outline the overarching task for later request execution context setup. Moving on to the next phase, they elaborate on extracting personas, entities, actions, and benefits from a story. Subsequently, they present an illustrative example of such extraction using a manually annotated story. Transitioning to the user role, they then supply the <STORY> to be processed. Given the statelessness of the model, the answer from the previous phase needs to be injected into the conversation. Consequently, adopting the assistant role, they introduce a conversation entry detailing the <CONCEPTS> acquired earlier. Following a similar pattern as in the preceding phase, they articulate the task using the system role (categorizing primary and secondary actions and entities) and provide an example of such an execution. The final step mirrors the previous patterns. They inject the <CATEGORIES> as the assistant, followed by task description and example provision in the system role. The conversation they used is described in Fig. 3.
Figure 3. An example of work process for extracting domain model
Zhang et al. (Zhang et al., [n. d.]) conducts a comprehensive evaluation of the performance of ChatGPT in the task of requirement information retrieval. This assessment offers profound insights for the design of more efficient retrieval methods. The evaluation results indicate that ChatGPT successfully retrieves relevant requirement information. However, under zero-shot conditions, its precision in retrieving specific requirement information is subject to certain limitations. These findings provide valuable guidance for further optimizing the application of ChatGPT in the field of requirement information retrieval.
Waseem et al. (Waseem et al., 2023) provides an in-depth exploration of the effects of students utilizing ChatGPT at different stages of the software development life cycle. Specifically, they provides students with a software project topic for development. During the requirements analysis phase, students gather functional and non-functional requirements with ChatGPT to understand project needs comprehensively. ChatGPT’s assistance is crucial for subsequent development stages. Upon completion of the development, Waseem et al.(Waseem et al., 2023) makes a survey to evaluate the effectiveness of ChatGPT, revealing a significant reduction in the time required for defining requirements during the requirements analysis phase. Additionally, students respond positively to the utility of ChatGPT, believing that its usage enhances requirement clarity. Also, ChatGPT aids in a better understanding of project needs and facilitates stakeholder communication. To conclude, ChatGPT demonstrated a significant positive impact during the requirements analysis phase of software development projects.
@@ 2.2.1. Datasets
Chen et.al. (Chen et al., 2021) utilized a purely handwritten collection of programming problems called HumanEval [^2] https://www.github.com/openai/human-eval. This dataset is based on the Python programming language and consists of 164 programming problems. Each programming problem in HumanEval consists of a function name, a natural language statement describing the function of the function, and unit test cases used to evaluate the generated code. The average number of unit test cases for each programming problem is 7.1. These programming problems cover many types, including language comprehension, reasoning, algorithms, and simple mathematics.
To assess the ability of LLM to generate concise Python code, MBPP [^3] https://github.com/google-research/google-research/tree/master/mbpp. (Austin et al., 2021) (The Mostly Basic Programming Problems) is proposed. The dataset includes 974 short Python programming problems, and the difficulty of each programming problem is set at the entry-level programmer level. Among these problems, 58% are mathematical calculation, 43% involve list processing, 19% require string operations, 9% process integer sequences, and 2% are centered on using other data structures. Similar to HumanEval, each programming problem sample consists of a function name, the natural language describing the target code, and three unit test cases used to evaluate the generated code.
@@
Wang et.al. (Wang et al., 2023a) proposed a dataset called MCoNaLa [^6] https://github.com/zorazrw/multilingual-conala. with different natural languages as prompts. They annotated a total of 896 natural language-code pairs in three languages: Spanish, Japanese, and Russian. Specifically, they engaged annotators fluent in both the respective natural language and Python. The annotators manually collected 341, 210, and 345 natural language-code pairs from the Spanish, Japanese, and Russian subforums of StackOverflow, respectively.
@@
To further investigate the performance of LLMs in generating more complex code, Du et.al. (Du et al., 2023) proposed a class-level code generation benchmark dataset. They manually build a benchmark for 100 class-level Python code generation tasks, named ClassEval [^7] https://github.com/FudanSELab/ClassEval.. Each coding task in ClassEval comprises a natural language description of the target class, a test suite used to validate the correctness of the generated code, and a specification solution serving as a reference implementation for the target class. Typically, LLMs generate code snippets based on the input description and use the provided test suite to verify correctness. They build test cases on two levels: method-level testing and class-level testing. To ensure that the generated code can be effectively checked according to the given test suite, they designed class-level and method-level descriptions for the target code. Method-level testing mainly involves independently calling each method to verify its correctness without calling any other method in the class. Class-level testing mainly evaluates correctness through sequential calls.
Although the LLM code generation model has achieved exciting performance, it is often fragile, as slight modifications to prompts can result in significantly different outputs. Most existing work on robustness in text or code tasks has focused on classification, lacking benchmarks for testing the robustness of code generation. Therefore, Wang et.al. (Wang et al., 2023d) proposes ReCode, a comprehensive benchmark for evaluating the robustness of code generation models. They tailored over 30 perturbations specifically for code, including document strings, function and variable names, code syntax, and code formatting. These perturbations were applied to the HumanEval and MBPP datasets, forming a robustness benchmark test dataset for code generation.